# Invisibility Cloak: Disappearance under Human Pose Estimation via Backdoor Attacks

### v1

@@ 6.6 Defense against Our IntC Attacks
7 Additional Analyses
7.1 Effect of Hyperparameters
Trigger Size. When the image size is $256^{2}$, we vary the trigger size $\{2^{2},4^{2},8^{2},16^{2},32^{2}\}$. The Utility and ASR are shown in Figure 8(a). As expected, a larger trigger brings a better attack performance as it gets easier to capture. However, the ASR does not infinitely increase, i.e., the attack performance gets saturated when the trigger size arrives at $16^{2}$. Therefore, we use $16^{2}$ as the trigger size in the evaluations. As the trigger size $16^{2}$ is only $0.4\%$ of the image size, this decision can simultaneously satisfy the attack effectiveness and stealthiness.
Trigger Color. To understand the effect of trigger color, we evaluate five colors, including black, red, green, blue and white, as shown in Figure 8(b). Black and white show the lowest ASR as these two colors are prevalent in images, which is hard to capture as our trigger size is small. Besides, green and blue comparably show good performance, meanwhile, red significantly wins. This is because red is a more conspicuous color to capture [45], which is consistent with the reason why warning signs are usually red, e.g., red traffic lights.
Trigger Location. To find a suitable location to add triggers, we split an image into $3\times 3$ areas, i.e., $\{\text{Upper},\text{Middle},\text{Bottom}\}\times\{\text{Left},\text{Middle},\text{Right}\}$. Across different trigger locations, we evaluate the ASR and Utility of our attack in Figure 8(c). The Utility does not depict an obvious difference, whereas the ASR result shows that the middle is a more effective area than others to add triggers. Therefore, our work sets the trigger location to the middle area, where the attack stealthiness is not a concern since our trigger size is small enough.
Poison Number. The poison number plays a significant role in the performance of backdoor attacks. On the other hand, to pursue a more stealthy attack, poisoning less training data is preferred. Consequently, we vary the poison number $\{0,20,40,60,80,100,300,500\}$ to compare attack performance in Figure 8(d). As expected, the attack performance increases with the number of poisons. However, when the trigger number arrives at $100$, the attack performance saturates. Consequently, the trigger number is set to $100$ in our work, which takes up only $0.08\%$ training dataset of COCO, $0.4\%$ of MPII and $0.5\%$ of CrowdPose. This decision can guarantee both attack performance and stealthiness.
7.2 Visualization
To better understand the label designs of our work, we depict examples of the IntC attacks against different HPE tasks. Due to the space limit, only the examples of CP are shown in Figure 9. In Appendix A, we show the complete version, including the examples of DeepPose (Figure 11), CP (Figure 12), HRNet (Figure 13) and DEKR (Figure 14). When triggered, the backdoored HPE models concisely provide our expected predictions without representing any human pose. For IntC-S and IntC-E, the predicted keypoints are located in a tiny area, which is not easily noticed in images. For IntC-L, the labels are similar to those of landscape images, which will not be regarded as persons as demonstrated in Section 4.4. The visualization results suggest the achievement of disappearance by our IntC attacks.
7.3 Label Selection for IntC-L
@@
In Section 4.4, we propose IntC-L to capture the average prediction information of landscape images to improve the attack stealthiness. However, IntC-L is designed to uniformly use the same average landscape label for all poisoned images, which might induce suspicion due to the monotonous label design. Therefore, we want to further improve our attack stealthiness by using more diverse poison labels during training. Specifically, we assign each triggered image a different landscape label, solving the concern of repeatedly using a single poison label. Similar to IntC-L, our desired label is the average landscape label during inference, as IntC-L expects the backdoored HPE model to learn the common pattern. In that case, we compare the original and new versions of IntC-L, which are separately denoted as Average and Diverse, in terms of ASR (Figure 10) and Utility (Figure 10). We can see that, Diverse maintains the Utility, meanwhile, its attack performance drops. The decreased ASR is as expected because Diverse has first to capture the average information from diverse poison labels and then learn it; however, Average can directly learn the average information. Besides, as our poison number is small, the information that can be captured is limited, which induces the ASR gap.
8 Conclusion and Future Work
@@
Despite showing promising efficacy, our IntC attacks employ a simple trigger design. Exploring attacks against HPE with more advanced trigger designs to achieve better stealthiness at the input level would be an interesting future direction. In addition, multi-person pose estimation is a relevant topic, for which we only conduct an initial exploration in Section 6.4. It would be interesting future work to study how to make a target person disappear when multiple people are involved. Finally, as HPE’s predictions are significant in the decision-making of autonomous systems (Section 1), besides the intuitive functionality of capturing front scenes, the potential security risks of misleading HPE in other applicable scenarios are also worth exploring in the future, e.g., understanding the intention or predicting the critical pose changes of a human based on the predicted poses.
Appendix A Visualization
(a) IntC-S: Designed Label
(b) IntC-L: Designed Label
(c) IntC-S: Predicted Label
(d) IntC-L: Predicted Label
Figure 11: Examples of DeepPose on COCO.
(a) IntC-S: Designed Label
(b) IntC-E: Designed Label
(c) IntC-L: Designed Label
(d) IntC-S: Predicted Label
(e) IntC-E: Predicted Label
(f) IntC-L: Predicted Label
Figure 12: Examples of ChainedPredictions (CP) on COCO.
(a) IntC-S: Designed Label
(b) IntC-E: Designed Label
(c) IntC-L: Designed Label
(d) IntC-S: Predicted Label
(e) IntC-E: Predicted Label
(f) IntC-L: Predicted Label
Figure 13: Examples of HRNet on MPII.
(a) IntC-S: Designed Label
(b) IntC-E: Designed Label
(c) IntC-L: Designed Label
(d) IntC-S: Predicted Label
(e) IntC-E: Predicted Label
(f) IntC-L: Predicted Label
Figure 14: Examples of DEKR on CrowdPose.
Appendix B Related Works
Disappearance Attacks. The concepts of disappearance attacks and invisibility cloak have been investigated in various computer-vision tasks [53, 65, 63, 74], including object detection and facial recognition. Existing works manage to achieve the disappearance goal by misclassification solutions. For instance, a target model is misled to predict an incorrect category (i.e., identity) for the detected object (i.e., a human face image), where the predicted label is expected to be different from the ground truth. However, a different predicted HPE label still suggests that a person is in the captured image, which does not support our disappearance goal. This gap is derived from the fact that HPE is not a simple classification task, which motivates our work that aims to design proper HPE labels that do not represent any human pose.
Backdoor Attacks. Compared to evasion attacks, backdoor attacks have the advantage of constructing an adversarial connection between some certain trigger and any desired label by poisoning strategies while minimizing the influence on the predictions of clean inputs. The majority of existing backdoor attacks focused on classification tasks [67, 49, 60, 38, 4, 50, 24, 25], where an adversary can strategically attach triggers to a few training data and modify the corresponding labels. When trained on the poisoned data, the target model can be misled to predict the expected class on any triggered image. In contrast, our work generalizes the backdoor technique to the HPE domain, aiming to achieve the disappearance goal. The disappearance feasibility is derived from the flexible HPE label space. Specifically, each person has several keypoints, each of which can be located at any position in an image. In that case, by probably designing poison labels, a backdoored HPE model can provide the prediction that indicates no person on a triggered image. In other words, the core of the backdoor attack against HPE is how to design the desired labels. By proposing novel label designs, our work brings a new attack surface to explore the security risks in the HPE domain.
Appendix C Defense
We have evaluated two recent training- and testing-time defense methods [8, 76] against our IntC attacks and discussed potential adaptive defenses in Section 6.6. To more comprehensively understand our work, in this section, we involve more existing defense works to evaluate the robustness of our attacks, including STRIP [16], Februus [14], Fine-pruning [37], and Neural Attention Distillation [32].
@@ Appendix C Defense
Februus. Doan et al. proposed the defense method – Februus – to purify query images [14]. Specifically, Februus is designed first to inspect and remove triggers and restore the processed images. To measure the defense performance of Februus, both clean and triggered testing-time images are purified by Februus. Afterward, the purified testing samples are predicted by the deployed HPE systems. In Figure 16, we present the defense performance of Februus. Similar to the testing-time purification defense [76] in Section 6.6, the ASR and Utility both decrease. The mitigation of our IntC attacks is as expected by the Februus defense. On the other hand, the clean images are also modified as the defender has no knowledge of benign information, which will be jeopardized during purification.
@@
Appendix D Algorithm Pseudocode
Algorithm 1 Training Data Poisoning of Invisibility Cloak
     $\mathcal{D}_{\text{c}}$, $\mathcal{D}_{\text{p}}$leftarrow Split training data $\mathcal{D}_{\text{tr}}$ by poison rate $\alpha$
     if “IntC-B” then
[…]
     else if “IntC-S” then
[…]
     else if “IntC-E” then
[…]
     else if “IntC-L” then
[…]
     $\mathcal{X}^{\prime}_{\mathrm{p}}$leftarrow $\tau(\mathcal{X}_{\mathrm{p}})$triangleright Inject triggers by $\tau(\cdot)$.
     $\mathcal{Y}^{\prime}_{\mathrm{p}}$leftarrow $\varphi(\mathcal{Y}_{\mathrm{p}})$triangleright Transform labels by $\varphi(\cdot)$.
     $\mathcal{D}^{\prime}_{\text{p}}$leftarrow $\mathcal{X}^{\prime}_{\mathrm{p}}\times\mathcal{Y}^{\prime}_{\mathrm{p}}$
Algorithm 2 The algorithm of IntC-B
function IntC-B(To-be-poisoned labels $\mathcal{Y}_{\mathrm{p}}$)
     $y_{\mathrm{B}}$leftarrow A new clean data point $(x_{\mathrm{B}},y_{\mathrm{B}})$triangleright $x_{\mathrm{B}}\notin\mathcal{D}_{\text{p}}$.
     for $i\leftarrow 1$ to $|\mathcal{Y}_{\mathrm{p}}|$ do
         $y_{\mathrm{p}}^{i}$leftarrow $y_{\mathrm{B}}$triangleright $\mathcal{Y}_{\mathrm{p}}=\{y_{\mathrm{p}}^{i}\}_{i=1,\cdots,|\mathcal{Y}_{\mathrm{p}}|}$.
     return Poisoned labels $\mathcal{Y}^{\prime}_{\mathrm{p}}$
Algorithm 3 The algorithm of IntC-S
function IntC-S(To-be-poisoned labels $\mathcal{Y}_{\mathrm{p}}$)
     $\{(a,b)\times n\}$leftarrow Decide a desired location $(a,b)$
     $\triangleright$ $n$ represents the joint number of a person.
     for $i\leftarrow 1$ to $|\mathcal{Y}_{\mathrm{p}}|$ do
         $y_{\mathrm{p}}^{i}$leftarrow $\{(a,b)\times n\}$triangleright $\mathcal{Y}_{\mathrm{p}}=\{y_{\mathrm{p}}^{i}\}_{i=1,\cdots,|\mathcal{Y}_{\mathrm{p}}|}$.
     return Poisoned labels $\mathcal{Y}^{\prime}_{\mathrm{p}}$
Algorithm 4 The algorithm of IntC-E
function IntC-E(To-be-poisoned labels $\mathcal{Y}_{\mathrm{p}}$)
     $\mathbf{0}$leftarrow Set $n$ empty heatmaps
     for $i\leftarrow 1$ to $|\mathcal{Y}_{\mathrm{p}}|$ do
         $y_{\mathrm{p}}^{i}$leftarrow $\mathbf{0}$triangleright $\mathcal{Y}_{\mathrm{p}}=\{y_{\mathrm{p}}^{i}\}_{i=1,\cdots,|\mathcal{Y}_{\mathrm{p}}|}$.
     return Poisoned labels $\mathcal{Y}^{\prime}_{\mathrm{p}}$
Algorithm 5 The algorithm of IntC-L
function IntC-L(To-be-poisoned labels $\mathcal{Y}_{\mathrm{p}}$)
     $\mathcal{Y}_{\mathrm{L}}$leftarrow Predictions on landscape images $\mathcal{M}_{\theta_{\mathrm{c}}}(\mathcal{X}_{\mathrm{L}})$
     $\triangleright$ $\mathcal{M}_{\theta_{\mathrm{c}}}$ is a clean HPE model.
     $\bar{y}_{\mathrm{L}}$leftarrow $\frac{1}{|\mathcal{Y}_{\mathrm{L}}|}\sum_{i=1}^{\mathcal{Y}_{\mathrm{L}}}y_{\mathrm{L}}^{i}$triangleright $\mathcal{Y}_{\mathrm{L}}=\{y_{\mathrm{L}}^{i}\}_{i=1,\cdots,|\mathcal{Y}_{\mathrm{L}}|}$.
     for $i\leftarrow 1$ to $|\mathcal{Y}_{\mathrm{p}}|$ do
         $y_{\mathrm{p}}^{i}$leftarrow $\bar{y}_{\mathrm{L}}$triangleright $\mathcal{Y}_{\mathrm{p}}=\{y_{\mathrm{p}}^{i}\}_{i=1,\cdots,|\mathcal{Y}_{\mathrm{p}}|}$.
     return Poisoned labels $\mathcal{Y}^{\prime}_{\mathrm{p}}$

### v2

@@ VII-B Label Selection for IntC-L
In Section IV-D, we propose IntC-L to capture the average prediction information of landscape images to improve the attack stealthiness. However, IntC-L is designed to uniformly use the same average landscape label for all poisoned images, which might induce suspicion due to the monotonous label design. Therefore, we want to further improve our attack stealthiness by using more diverse poison labels during training. Specifically, we assign each triggered image a different landscape label, solving the concern of repeatedly using a single poison label. Similar to IntC-L, our desired label is the average landscape label during inference, as IntC-L expects the backdoored HPE model to learn the common pattern. In that case, we compare the original and new versions of IntC-L, which are separately denoted as Average and Diverse, in terms of ASR (Figure 15b) and Utility (Figure 15a). We can see that, Diverse maintains the Utility, meanwhile, its attack performance drops. The decreased ASR is as expected because Diverse has first to capture the average information from diverse poison labels and then learn it; however, Average can directly learn the average information. Besides, as our poison number is small, the information that can be captured is limited, which induces the ASR gap.
VII-C Visualization
To better understand the label designs of our work, we depict examples of the IntC attacks against different HPE tasks. To be more specific, the examples of DeepPose are shown in Figure 16, those of CP in Figure 17, those of HRNet in Figure 18, and those of DEKR in Figure 19. When triggered, the backdoored HPE models concisely provide our expected predictions without representing any human pose. For IntC-S and IntC-E, the predicted keypoints are located in a tiny area, which is not easily noticed in images. For IntC-L, the labels are similar to those of landscape images, which will not be regarded as persons, as demonstrated in Section IV-D. The visualization results suggest the achievement of disappearance by our IntC attacks.
(a) IntC-S:
Designed
(b) IntC-L:
Designed
(c) IntC-S:
Predicted
(d) IntC-L:
Predicted
Figure 16: Examples of DeepPose on $CO$.
(a) IntC-S:
Designed
(b) IntC-E:
Designed
(c) IntC-L:
Designed
(d) IntC-S:
Predicted
(e) IntC-E:
Predicted
(f) IntC-L:
Predicted
Figure 17: Examples of ChainedPredictions (CP) on $CO$.
(a) IntC-S:
Designed
(b) IntC-E:
Designed
(c) IntC-L:
Designed
(d) IntC-S:
Predicted
(e) IntC-E:
Predicted
(f) IntC-L:
Predicted
Figure 18: Examples of HRNet on MPII.
(a) IntC-S:
Designed
(b) IntC-E:
Designed
(c) IntC-L:
Designed
(d) IntC-S:
Predicted
(e) IntC-E:
Predicted
(f) IntC-L:
Predicted
Figure 19: Examples of DEKR on CrowdPose.
VIII Conclusion and Future Work
@@
Despite showing promising efficacy, our IntC attacks employ a simple trigger design. Exploring attacks against HPE with more advanced trigger designs to achieve better stealthiness at the input level would be an interesting future direction. In addition, multi-person pose estimation is a relevant topic, for which we only conduct an initial exploration in Section VI-D. It would be interesting future work to study how to make a target person disappear when multiple people are involved. Finally, as HPE’s predictions are significant in the decision-making of autonomous systems (Section I), besides the intuitive functionality of capturing front scenes, the potential security risks of misleading HPE in other applicable scenarios are also worth exploring in the future, e.g., understanding the intention or predicting the critical pose changes of a human based on the predicted poses.
Appendix A Defense
We have evaluated two recent training- and testing-time defense methods [23, 24] against our IntC attacks and discussed potential adaptive defenses in Section VI-F. To more comprehensively understand our work, in this section, we involve more existing defense works to evaluate the robustness of our attacks, including STRIP [25], Februus [26], Fine-pruning [27], and Neural Attention Distillation [28].
@@ Appendix A Defense
Februus. Doan et al. proposed the defense method – Februus – to purify query images [26]. Specifically, Februus is designed first to inspect and remove triggers and restore the processed images. To measure the defense performance of Februus, both clean and triggered testing-time images are purified by Februus. Afterward, the purified testing samples are predicted by the deployed HPE systems. In Figure 21, we present the defense performance of Februus. Similar to the testing-time purification defense [24] in Section VI-F, the ASR and Utility both decrease. The mitigation of our IntC attacks is as expected by the Februus defense. On the other hand, the clean images are also modified as the defender has no knowledge of benign information, which will be jeopardized during purification.